Assume stop-and-wait, 100 Mbps link, 10 ms propagation delay, 10,000 bit packet. Calculate sender utilization. Assume...

  1. Assume stop-and-wait, 100 Mbps link, 10 ms propagation delay, 10,000 bit packet. Calculate sender utilization.
  1. Assume pipelining and 20 packets can be sent without ACK, 100 Mbps link, 10 ms propagation delay, 10,000 bit packet. Calculate sender utilization.
Answer #1

a.

Given that bandwidth = 100 Mbps = 100 * 10^6 bits/sec

Length of the packet = 10, 000 bits

So, transmission delay(Tt) = length / bandwidth = 10000/(100* 106) = 0.0001 sec = 0.1 ms

and propagation delay(Tp) is given as 10 ms.

So, sender utilization / efficiency = Tt / (Tt + 2*Tp) = 0.1 / ( 0.1 + 2*10) = 0.1 / (20.1) = 0.499 %

b.

Transmission delay (Tt) = length / bandwidth = 10000/(100* 106) = 0.0001 sec = 0.1 ms

Propagation delay ( Tp) = 10 ms.

And number of packets(N)  at a time given is 20.

Then a is (Tp/Tt) = 100

So, sender utilization / efficiency = N / ( 1 + 2a) = 20 / ( 1 + 2*100) = 20/ 201 = 9.95 % efficiency